To ensure a growing season of sufficient length, Mang Popoy has at most 16 days left to




plant corn and soybean . he can plant com at a rate of 10 hectares per day and soybean at




15 hectares per day. If he has at most 200 hectares available, how many hectares of each




type of crop can he plant?​

1
Expert's answer
2022-03-11T06:33:45-0500

Consider the system of equations:

"\begin{cases}
10a+15b=200,\\
a,b\geq0,\\
a+b\leq16;
\end{cases}"

Where "a" is the number of days when corn is planted, "b" – soybean.

From first equation:

"a=20-1.5b;"

Since a farmer can plant one crop per day, then "b" must be even, otherwise "a" will not be an integer. Let's iterate over all the values of "b\in\{0..16,\ 2|b\}" as long as two inequalities from the system are valid:

Answer: "(a,b)=\{(8,8),(5,10),(2,12)\}"